OpenStudy (anonymous):

Your furniture store sells two types of dining room tables. The first, type A, costs $265 and you make a $25 profit on each one. The second, type B, costs $100 and you make a $13 profit on each one. You can order no more than 40 tables this month, and you need to make at least $760 profit on them. If you must order at least one of each type of table, how many of each type of table should you order if you want to minimize your cost? 20 of type A; 20 of type B 2 of type A; 38 of type B 30 of type A; and 10 of type B 38 of type A; 2 of type B

OpenStudy (anonymous):

a+b ≤ 20 C(a,b)=265a+100b 25a+13b ≥ 760 a+b ≤ 20 25a+13b ≥ 760 25a+25b ≤ 500 25a+13b ≥ 760 25a ≤ 500 - 25b 25a ≥ 760 - 13b 25a ≤ 25a 760 - 13b ≤ 500 - 25b 25b - 13b ≤ 500 - 760 12b ≤ -260 b ≤ -260/12 b ≤ -21.6

OpenStudy (anonymous):

A + B ≤ 40

OpenStudy (anonymous):

How did you get that ?

OpenStudy (anonymous):

" You can order no more than 40 tables this month,"

OpenStudy (anonymous):

is this rite for the next step ?

OpenStudy (anonymous):

C(a,b)=265a+100b C(40,-21.6)=265(40)+100(-21.6)=8440

OpenStudy (anonymous):

ok . . .

OpenStudy (anonymous):

A + B ≤ 40 -> 25 A + 13 ( 40 - A) ≥ 760 => A ≥ ( 760 - 520 ) / 12 = 20 tables

OpenStudy (anonymous):

wow , im sooo confused .

OpenStudy (anonymous):

-> B ≤ 20 tables Thus C ( A, B) = 265* 20 + 100 * 20 = $ 7,300

OpenStudy (anonymous):

So , its the 1st answer ?

OpenStudy (anonymous):

All I see is one question, minimize cost?

OpenStudy (anonymous):

What ?

OpenStudy (anonymous):

" how many of each type of table should you order if you want to minimize your cost?

OpenStudy (anonymous):

Yeah , and is it the first answer ? 20 of type A; 20 of type B

OpenStudy (anonymous):

I guess that's all !

OpenStudy (anonymous):

Well , thank you . you helped alot ! I appreciate it (:

OpenStudy (anonymous):

Did you see how I solve the equation, make it short it'll be easier to double check!

OpenStudy (anonymous):

Yeah , pretty much ! (: